A 117-cm-long, 5.00 g string oscillates in its n = 2 mode with a frequency of 200 Hz and a maximum amplitude of 4.50 mm .

Part A:

What is the wavelength in the string?

Express your answer with the appropriate units.

Part B:

What is the tension in the string?

Express your answer with the appropriate units.
Answer #1

A)

Wavelength, lambda = L/2 = 1.17/2 = 0.585 m

B)

200 = (2/ 2 x 1.17) x sqrt(T/(0.005/1.17))

Tension, T = 234 N

Comment in case any doubt please rate my answer....